# Pre-Training Representations of Binary Code Using Contrastive Learning

### v4

@@ Abstract
Index Terms:
Technical articles in SE, artificial intelligence, binary code analysis, contrastive learning
@@ 1 Introduction
Bibary code [^1] In this paper, we use binary code to refer to both executable binary functions and lifted assembly/intermediate representations because they are trivially interchangeable. provides valuable information about a program’s content and behavior and is often the only available representation of a program in certain cases, such as legacy systems, proprietary software, and penetration testing [1]. Binary code analysis is critical to many software related tasks, and understanding binary code can serve as a critical source of information about program content and behavior, and thus a foundation of many applications [2, 3, 4, 5].
@@
Currently, binary code analysis is approached through two methods: (1) traditional static and dynamic analysis and (2) AI-based methods. Traditional methods use manual techniques and specific algorithms (e.g., dataflow analysis), but are limited in their ability to be used across different platforms and applications [6, 7, 8]. AI-based methods use machine learning to capture both program structure and semantics, and have combined syntax, semantics, and structure to create embeddings for specific tasks [9, 10, 11]. However, few focus on large-scale pre-trained representations for binary code, which is a largely unexplored area [12, 9], instead focusing primarily on source code or related syntactic structures like abstract syntax trees.
@@
The design of ContraBin is based on three key insights derived from multi-view learning, human translation and code compilation, and contrastive learning: (1) source code, corresponding human-written comments, and compiled binary code can flexibly represent the same semantics of a given program but just in different modalities. (2) Human documentation (e.g., comments within source code, written in natural languages) and compilation (or decompilation) by the compiler share similar gradual learning properties, as shown in Figure 1. With this insight, we design simplex interpolation among source code, binary code, and comments to imitate a unified learning process. (3) While source code, binary code, comments, and their intermediate representations correspond to the same program semantics, they are dissimilar to other source code, binary code, comments, intermediate representations. We show how these insights can be combined to form an effective multi-modal contrastive learning embedding capable of aiding the analysis of binaries.
We evaluate ContraBin’s pre-trained embeddings against state-of-the-art large-scale pre-trained models [13, 12, 9, 14] [^2] We note that existing pre-trained models have focused on applications where source code is available — we make a best effort to compare our novel embedding for binaries against previous work in spite of this key difference.. We evaluate ContraBin in four indicative downstream tasks and corresponding datasets that entail binary code analysis: (1) binary functional algorithm analysis, (2) binary function name recovery, (3) binary code summarization, and (4) binary reverse engineering, using POJ-104 [15], DIRE [10], and two subsets of AnghaBench [16], respectively.We choose these tasks and datasets because they also permit a fair comparison with state-of-the-art models that make use of source code. Our results show that ContraBin substantially outperforms current code analysis models on the vast majority of task-relevant metrics and achieves comparable performance on others. In conclusion, we claim the following contributions:
@@
    We present ContraBin, a novel contrastive learning-based framework leveraging simplex interpolation on source code, comments, and binary code to learn an effective representation (i.e., a pre-trained embedding) of binary code.
@@
    We design a new simplex interpolation approach to capture the inherent process of gradual learning of humans to assist contrastive learning.
@@
    We derive a novel intermediate contrastive learning algorithm to incorporate external knowledge and semantics in source code and comments into binary code.
@@
    We provide an in-depth analysis of the impact of human-written and synthetic comments on binary code comprehension tasks, demonstrating that synthetic comments significantly enhance performance, while human-written comments can introduce noise and reduce model effectiveness compared to using no comments.
  • $\bullet$
    To the best of our knowledge, this is the first study to derive intermediate contrastive learning for binary code analysis to improve representation model performance.
  • $\bullet$
    We conduct extensive experiments on indicative datasets in evaluation, demonstrating that ContraBin outperforms other state-of-the-art models in all four downstream tasks on binary code analysis with respect to task-relevant metrics such as BLEU-4 for summarization and accuracy for algorithmic classification.
@@ 2.1 Primary Contrastive Learning
This subsection introduces the primary contrastive learning approach for the simplex interpolation. Recall that the conversion process between two representations can involve complex analyses (e.g., compilation) or creative human processes (e.g., code comprehension). At the same time, no generalizable models have been trained on datasets containing binary code. This can lead to the cold start problem in representation learning [17] when a new data representation (e.g., a novel program never seen before) emerges in a system. Therefore, it is challenging to capture their features — the same scenario applies to binary code representation.
@@
To mitigate this cold start problem, we adopt primary contrastive learning in the first step of model training. Specifically, we use three representations of a program: comments (i.e., summarization generated by code models), source code (i.e., as written by developers), and binary code (i.e., as emitted from a compiler), and randomly choose two of them to conduct contrastive learning, as shown in Figure 2. During this step, the simplex interpolation module is disabled: this approach is inspired by multi-view learning [18] in computer vision, in which an object can have $k$ views, and all of the views are exactly from the same object. In our model, either source code, binary code, or comment can represent the same program, so we treat the three representations (also denoted as modalities or views) of a program to enrich the information in binary code There are two steps in primary contrastive learning: manifold projection and batch-wise similarity comparison.
@@
Batch-wise similarity comparison To align binary code representations with their corresponding source code and comments, we adopt a loss function inspired by CLIP [19]. This method leverages batch-wise similarity and dissimilarity between different program representations to update the model.
@@ 2.2 Secondary Contrastive Learning
We use both linear and non-linear simplex interpolation as part of a gradual learning approach to obtain increasingly expressive embeddings. Based on simplex interpolation theories [20], we presume the intermediate representations between any of the two in source code, binary code, and comments can be a different view of the same program. The detailed steps are shown as follows:
@@
Gradual learning Following the gradual learning literature, models learn most effectively when they follow a natural, progressive order as humans do [21, 22]. Specifically, models learn high-level data abstracts initially, then progress to more complicated embeddings. Following this intuition, we train our model via primary contrastive learning during cold start. After several epochs (10 epochs in our implementation), we switch the training pattern to contrastive learning based on linear simplex interpolation and non-linear simplex interpolation to further improve the generalizability of our model and finalize the training process, shown in Algorithm 1. By integrating all three parts, ContraBin follows a natural learning curve (i.e., from simple to complex) to better understand the meaning of binary code.
@@
0:  Source code set $X_{s}$ with paired binary code set $X_{b}$ and comment set $X_{c}$, split into train $X^{train}$ and validation $X^{val}$.
0:  Minimization of loss on $X^{val}$.
1:  if Primary contrastive learning then
2:     for $batch=1,\ldots,k_{start}$ do
3:        Project $X_{c}$, $X_{s}$ and $X_{b}$ to $\mathcal{M}$ with encoders $f_{a}$ and $f_{t}$
4:        Compute batch-wise similarity loss between two representations in $\mathcal{M}$ and train the predictors and encoder
5:     end for
6:  end if
7:  if Contrastive learning by linear or non-linear simplex interpolation then
8:     for $batch=1,\ldots,k_{interp}$ do
9:        Project $X_{c}$, $X_{s}$ and $X_{b}$ to $\mathcal{M}$ with encoder $f_{a}$ and $f_{t}$
10:        Choose two of the encoded representations and compute the index $\lambda_{l}$ or $\lambda_{n}$ by the interpolation NN and get intermediate representations
11:        Compute batch-wise similarity loss between intermediate and the other representation in $\mathcal{M}$ and train the binary encoder and interp NN
12:     end for
13:  end if
@@ 2.3 Task-Specific Fine-Tuning
TABLE I: Categorization and objectives of downstream tasks.
@@
Algorithmic Functionality Classification from Binaries This task involves classifying binary code according to its functionality, such as distinguishing ”quicksort” from ”md5 hash.” Accurate classification assists developers in isolating specific code blocks, improving readability, saving development time, and aiding in reverse engineering [23]. For example, given a binary, the goal is to categorize it into classes like sorting, hashing, or searching. As shown in Table I, this task is categorized under ”Analysis” and represents a key aspect of binary code analysis, where our method leverages multiple program representations for robust classification.
Function Name Recovery Recovering function names from binary code sequences is crucial for stripped binaries that lack meaningful names and debugging information. This task, categorized as ”Analysis” in Table I, significantly reduces manual effort in reverse engineering and malware analysis scenarios, particularly when source code is unavailable. By predicting function names effectively, our embeddings restore clarity to binary code, enhancing the overall analysis workflow.
Code Summarization This task translates binary code into concise English summaries, aiding comprehension and facilitating security assessments, reverse engineering, and software maintenance. As outlined in Table I, code summarization falls under the ”Comprehension” category, emphasizing its role in improving understanding of binary functionality and enabling more effective software analysis through succinct summaries.
Reverse Engineering Decompiling binary code into source code is essential for tasks like security assessments, defect localization, and software comprehension. While tools like Hex-Rays and Ghidra offer decompilation support, they often generate code that lacks symbolic names and deviates from canonical developer-written formats. Reverse engineering, classified under ”Comprehension” in Table I, benefits from our embeddings by transforming binaries into more interpretable source code, providing valuable insights into the original binaries.
We summarize the categorization of these tasks, their objectives, and their respective focus areas in Table I, which groups the tasks into ”Analysis” and ”Comprehension” domains.
@@ 3.1 Training Design
Source code: To obtain source code, we use a widely-used public dataset called AnghaBench [16] as our source data for training ContraBin [^4] AnghaBench: http://cuda.dcc.ufmg.br/angha/home. Specifically, we adopt the benchmark dataset that contains 1 million single-function C files extracted from programs and mined from popular GitHub repositories. This dataset serves as the source code representation in our method to incorporate semantic variety into the training process.
@@
Comments: During pre-analysis, we found most comments in real-world source code are not globally informative. Specifically, human-written comments can be random in content (e.g., they may contain copyright notifications, random snippets of old code, or unstructured explanations), or contain partial information about the current location of code statements — source code comments in general are not always descriptions of semantics of a program’s complete source code. Therefore, we adopt an Encoder-Decoder CodeT5 [24] model to automatically generate a single comment for each snippet of source code in our dataset.
@@
We use two Nvidia A40 GPUs during model pre-training and follow the parameter settings of SimpleCLIP [25]. We set the random seed as 42 in the pre-training for reproduction. To better evaluate the performance and robustness, we also train two versions of our model, ContraBin-PCL and ContraBin. For ContraBin-PCL, we set 10 epochs for primary contrastive learning only. For ContraBin, we use 10, 10, and 10 to improve the overall model efficiency on binary code analysis, and 10, 30, 30 to enhance the general model performance on binary code comprehension.
@@ 3.2 Evaluation Design
POJ-104 For the binary functional algorithm classification task (RQ1: Downstream Task 1), we adopt POJ-104 from the CodeXGLUE dataset [26]. In an Open Judge (OJ) system, students submit their solutions to a programming problem, and the OJ system judges whether the code can run successfully on all available test suites. In this way, the OJ system aims to improve the programming skills of users. From this dataset, we can find other programs that perform the same task as an input program (e.g., programs can be classified as bubble sort vs. heap sort. vs. Fibonacci sequence, etc.). The dataset includes programming problems and verified source code solutions. There are 104 programming problems in POJ-104, with 500 examples each. The dataset is categorized into train/development/test sets with non-overlapping program problem labels. Thus, we can view this dataset as containing 104 different classes in which to classify an input program. While the original POJ-104 dataset used source code structural information, we adapt it for classifying input binaries.
@@
In our experiments, we compile each solution in POJ-104 to LLVM assembly code, keeping the functionality label unchanged, as shown in Table II. This change transforms POJ-104 into an assembly code dataset suiting our binary analysis interests. We fine-tune each model with 2 epochs and a block size of 400. We use training and validation batch sizes of 32 and 8, respectively, and choose learning rate to be 2e-5 and maximal gradient normalization to be 1. In all fine-tuning processes, we use the default random seed of 123456.
@@
DIRT For the binary function name recovery task (RQ2), we chose the DIRT dataset used to train DIRTY[10], a recent machine learning model augmenting decompiler outputs with variable type and name predictions. The full dataset contains around 1 million functions spread across 75,656 binaries that were mined from public GitHub repositories using GHCC [^6] GHCC: https://github.com/huzecong/ghcc. With the help of a decompiler, each binary is lifted into equivalent C pseudocode and processed into a list of lexemes and an implementation-defined format for enumerating source-level variable information. Because the DIRT dataset is a preprocessed dataset that does not conform with ContraBin’s expected inputs, we instead obtained the list of GitHub repositories that were used to construct the DIRT dataset and compiled each project ourselves. We modified GHCC to save the intermediate LLVM bitcode objects produced during compilation and disassemble each bitcode object into readable LLVM assembly using the LLVM disassembler.
@@
In our experiments, we select function names with number of functions larger than 200, and remove some function names that are not specific to a type of function (i.e., main, _, and __list_add) to make sure the model is trained on meaningful data. We further strip all LLVM files of their original function names (i.e., @TESTFUN0 as the name of the first function in an LLVM file) and measure ContraBin’s ability to classify these function names. The dataset statistics are shown in Table II. We fine-tune each model with 5 epochs and a block size of 256. We use training and validation batch sizes of 8 and 16, respectively, and choose learning rate to be 2e-5 and maximal gradient normalization to be 1. In all fine-tuning processes, we use the default random seed as 123456.
@@
AnghaBench For the binary code summarization and reverse engineering tasks, we use the AnghaBench test set [16] during pre-training, which includes code never been seen by the model. Specifically, we only select the main function in LLVM code and its paired source code for fine-tuning the code summarization and reverse engineering tasks. To better fit the capacity of pre-trained language models, we further truncate the length of binary code by only selecting the first 512 instructions. The dataset statistics are displayed in Table II. In this paper, we evaluate the model’s ability to translate and summarize binary code using the validation set and analyze its embedding using the test set.
TABLE II: Dataset statistics for algorithmic classification.

### v5

@@ 1 Introduction
Binary code [^1] In this paper, we use binary code to refer to both executable binary functions and lifted assembly/intermediate representations because they are trivially interchangeable. provides valuable information about a program’s content and behavior and is often the only available representation of a program in certain cases, such as legacy systems, proprietary software, and penetration testing (Harris & Miller, 2005). Binary code analysis is critical to many software related tasks, and understanding binary code can serve as a critical source of information about program content and behavior, and thus a foundation of many applications (Wang et al., 2000; Pierce & Mudge, 1994; Cifuentes et al., 1999; Giffin et al., 2004).
@@
Currently, binary code analysis is approached through two methods: (1) traditional static and dynamic analysis and (2) AI-based methods. Traditional methods use manual techniques and specific algorithms (e.g., dataflow analysis), but are limited in their ability to be used across different platforms and applications (Bao et al., 2014; David et al., 2016; Aslanyan et al., 2020). AI-based methods use machine learning to capture both program structure and semantics, and have combined syntax, semantics, and structure to create embeddings for specific tasks (Guo et al., 2020; Lacomis et al., 2019; Chen et al., 2022b). However, few focus on large-scale pre-trained representations for binary code, which is a largely unexplored area (Feng et al., 2020; Guo et al., 2020), instead focusing primarily on source code or related syntactic structures like abstract syntax trees.
@@
The design of ContraBin is based on three key insights. First, we recognize that source code, comments, and their compiled binaries are different but semantically equivalent representations of a program, forming a rich multi-modal view. Second, we observe that the process of a human documenting code and a compiler translating it share properties of gradual learning, as illustrated in Figure 1. We model this with simplex interpolation to create a unified learning process that bridges these different views. Finally, our approach leverages the principle of semantic dissimilarity: while these related artifacts represent the same program, they are inherently distinct from the artifacts of any other program. These insights form the foundation of our multi-modal contrastive learning framework, designed to produce effective embeddings for binary analysis.
To validate our approach, we evaluate ContraBin’s pre-trained embeddings against several state-of-the-art large-scale models (Liu et al., 2019; Feng et al., 2020; Guo et al., 2020; 2022a). While many existing models are designed primarily for source code, we make a best-effort comparison to establish a strong baseline for binary-focused representation learning. Our evaluation spans four indicative downstream tasks: (1) Binary Functional Algorithm Analysis on the POJ-104 dataset (Mou et al., 2016); (2) Binary Function Name Recovery on the DIRE dataset (Lacomis et al., 2019); and (3) Binary Code Summarization and (4) Binary Reverse Engineering, both on subsets of AnghaBench (Da Silva et al., 2021).
These tasks were chosen as they provide a comprehensive test of binary analysis and comprehension capabilities and permit a fair comparison with source-code-aware models. Our results show that ContraBin substantially outperforms current code analysis models on the vast majority of task-relevant metrics and achieves comparable performance on others.
In conclusion, we claim the following contributions:
@@
    We present ContraBin, the first contrastive learning framework, to our knowledge, for learning representations of LLVM Intermediate Representation (IR). It uniquely aligns three modalities (source code, IR, and comments) using a two-stage process with simplex interpolation to create meaningful intermediate views.
@@
    We design a novel simplex interpolation method, inspired by multi-view and curriculum learning, that models a gradual learning process to better align the high-level semantics of source code and comments with the low-level structure of binary code.
@@
    We provide a key empirical finding on training data quality: concise, LLM-generated comments significantly improve performance, while noisy, human-written comments degrade it, reshaping how natural language annotations should be leveraged in binary analysis.
@@
    Our extensive evaluation on four diverse tasks—functionality classification, name recovery, summarization, and reverse engineering—shows that ContraBin consistently and significantly outperforms strong pre-trained baselines, demonstrating the broad applicability of its embeddings.
Collectively, these contributions are embodied in ContraBin, a framework designed to significantly advance the state of binary code representation learning. By bridging the semantic gap between source code, natural language, and low-level binaries, our work provides a robust foundation for a wide range of security and reverse engineering applications. To facilitate further research and ensure the reproducibility of our results, we have made our complete implementation, pre-trained models, and all datasets publicly available. [^2] https://zenodo.org/records/15219264
@@ 2.1 Primary Contrastive Learning
This subsection introduces the primary contrastive learning approach for the simplex interpolation. Recall that the conversion process between two representations can involve complex analyses (e.g., compilation) or creative human processes (e.g., code comprehension). At the same time, no generalizable models have been trained on datasets containing binary code. This can lead to the cold start problem in representation learning (Contardo et al., 2014) when a new data representation (e.g., a novel program never seen before) emerges in a system. Therefore, it is challenging to capture their features — the same scenario applies to binary code representation.
@@
To mitigate this cold start problem, we adopt primary contrastive learning in the first step of model training. Specifically, we use three representations of a program: comments (i.e., summarization generated by code models), source code (i.e., as written by developers), and binary code (i.e., as emitted from a compiler), and randomly choose two of them to conduct contrastive learning, as shown in Figure 2. During this step, the simplex interpolation module is disabled:
this approach is inspired by multi-view learning (Zhao et al., 2017) in computer vision, in which an object can have $k$ views, and all of the views are exactly from the same object. In our model, either source code, binary code, or comment can represent the same program, so we treat the three representations (also denoted as modalities or views) of a program to enrich the information in binary code
There are two steps in primary contrastive learning: manifold projection and batch-wise similarity comparison.
@@
Batch-wise similarity comparison To align binary code representations with their corresponding source code and comments, we adopt a loss function inspired by CLIP (Radford et al., 2021). This method leverages batch-wise similarity and dissimilarity between different program representations to update the model.
@@ 2.2 Secondary Contrastive Learning
We use both linear and non-linear simplex interpolation as part of a gradual learning approach to obtain increasingly expressive embeddings. Based on simplex interpolation theories (Thrampoulidis et al., 2022), we presume the intermediate representations between any of the two in source code, binary code, and comments can be a different view of the same program. The detailed steps are shown as follows:
@@
Gradual learning Following the gradual learning literature, models learn most effectively when they follow a natural, progressive order as humans do (Gou et al., 2021; Cho & Hariharan, 2019). Specifically, models learn high-level data abstracts initially, then progress to more complicated embeddings. Following this intuition, we train our model via primary contrastive learning during cold start. After several epochs (10 epochs in our implementation), we switch the training pattern to contrastive learning based on linear simplex interpolation and non-linear simplex interpolation to further improve the generalizability of our model and finalize the training process, shown in Algorithm 1. By integrating all three parts, ContraBin follows a natural learning curve (i.e., from simple to complex) to better understand the meaning of binary code.
@@
1:Source code set $X_{s}$ with paired binary code set $X_{b}$ and comment set $X_{c}$, split into train $X^{train}$ and validation $X^{val}$
2:Minimization of loss on $X^{val}$
3:if Primary contrastive learning then
4:  for $batch=1,\ldots,k_{start}$ do
5:    Project $X_{c}$, $X_{s}$ and $X_{b}$ to $\mathcal{M}$ with encoders $f_{a}$ and $f_{t}$
6:    Compute batch-wise similarity loss between two representations in $\mathcal{M}$ and train predictors and encoder
7:  end for
8:end if
9:if Contrastive learning by linear or non-linear simplex interpolation then
10:  for $batch=1,\ldots,k_{interp}$ do
11:    Project $X_{c}$, $X_{s}$ and $X_{b}$ to $\mathcal{M}$ with encoders $f_{a}$ and $f_{t}$
12:    Choose two encoded representations and compute $\lambda_{l}$ or $\lambda_{n}$ using the interpolation $N$ to get intermediate representations
13:    Compute batch-wise similarity loss between intermediate and the other representation in $\mathcal{M}$ and train the binary encoder and interp NN
14:  end for
15:end if
@@ 2.3 Task-Specific Fine-Tuning
Table 1: Categorization and objectives of downstream tasks.
@@
Algorithmic Functionality Classification from Binaries This task involves classifying binary code according to its functionality, such as distinguishing "quicksort" from "md5 hash." Accurate classification assists developers in isolating specific code blocks, improving readability, saving development time, and aiding in reverse engineering (Haq & Caballero, 2021). For example, given a binary, the goal is to categorize it into classes like sorting, hashing, or searching. As shown in Table 1, this task is categorized under "Analysis" and represents a key aspect of binary code analysis, where our method leverages multiple program representations for robust classification.
Function Name Recovery Recovering function names from binary code sequences is crucial for stripped binaries that lack meaningful names and debugging information. This task, categorized as "Analysis" in Table 1, significantly reduces manual effort in reverse engineering and malware analysis scenarios, particularly when source code is unavailable. By predicting function names effectively, our embeddings restore clarity to binary code, enhancing the overall analysis workflow.
Code Summarization This task translates binary code into concise English summaries, aiding comprehension and facilitating security assessments, reverse engineering, and software maintenance. As outlined in Table 1, code summarization falls under the "Comprehension" category, emphasizing its role in improving understanding of binary functionality and enabling more effective software analysis through succinct summaries.
Reverse Engineering Decompiling binary code into source code is essential for tasks like security assessments, defect localization, and software comprehension. While tools like Hex-Rays and Ghidra offer decompilation support, they often generate code that lacks symbolic names and deviates from canonical developer-written formats. Reverse engineering, classified under "Comprehension" in Table 1, benefits from our embeddings by transforming binaries into more interpretable source code, providing valuable insights into the original binaries.
We summarize the categorization of these tasks, their objectives, and their respective focus areas in Table 1, which groups the tasks into "Analysis" and "Comprehension" domains.
@@ 3.1 Training Design
Source code: To obtain source code, we use a widely-used public dataset called AnghaBench (Da Silva et al., 2021) as our source data for training ContraBin [^4] AnghaBench: http://cuda.dcc.ufmg.br/angha/home. Specifically, we adopt the benchmark dataset that contains 1 million single-function C files extracted from programs and mined from popular GitHub repositories. This dataset serves as the source code representation in our method to incorporate semantic variety into the training process.
@@
Comments: During pre-analysis, we found most comments in real-world source code are not globally informative. Specifically, human-written comments can be random in content (e.g., they may contain copyright notifications, random snippets of old code, or unstructured explanations), or contain partial information about the current location of code statements — source code comments in general are not always descriptions of semantics of a program’s complete source code. Therefore, we adopt an Encoder-Decoder CodeT5 (Wang et al., 2021) model to automatically generate a single comment for each snippet of source code in our dataset.
@@
We use two Nvidia A40 GPUs during model pre-training and follow the parameter settings of SimpleCLIP (Shariatnia, 2021). We set the random seed as 42 in the pre-training for reproduction. To better evaluate the performance and robustness, we also train two versions of our model, ContraBin-PCL and ContraBin. For ContraBin-PCL, we set 10 epochs for primary contrastive learning only. For ContraBin, we use 10, 10, and 10 to improve the overall model efficiency on binary code analysis, and 10, 30, 30 to enhance the general model performance on binary code comprehension.
@@ 3.2 Evaluation Design
POJ-104 For the binary functional algorithm classification task (RQ1: Downstream Task 1), we adopt POJ-104 from the CodeXGLUE dataset (Lu et al., 2021). In an Open Judge (OJ) system, students submit their solutions to a programming problem, and the OJ system judges whether the code can run successfully on all available test suites. In this way, the OJ system aims to improve the programming skills of users. From this dataset, we can find other programs that perform the same task as an input program (e.g., programs can be classified as bubble sort vs. heap sort. vs. Fibonacci sequence, etc.).
The dataset includes programming problems and verified source code solutions. There are 104 programming problems in POJ-104, with 500 examples each. The dataset is categorized into train/development/test sets with non-overlapping program problem labels. Thus, we can view this dataset as containing 104 different classes in which to classify an input program. While the original POJ-104 dataset used source code structural information, we adapt it for classifying input binaries.
@@
In our experiments, we compile each solution in POJ-104 to LLVM assembly code, keeping the functionality label unchanged, as shown in Table 2. This change transforms POJ-104 into an assembly code dataset suiting our binary analysis interests. We fine-tune each model with 2 epochs and a block size of 400. We use training and validation batch sizes of 32 and 8, respectively, and choose learning rate to be 2e-5 and maximal gradient normalization to be 1. In all fine-tuning processes, we use the default random seed of 123456.
@@
DIRE For the binary function name recovery task (RQ2), we chose the DIRE dataset used to train DIRTY model(Lacomis et al., 2019), a recent machine learning model augmenting decompiler outputs with variable type and name predictions. The full dataset contains around 1 million functions spread across 75,656 binaries that were mined from public GitHub repositories using GHCC [^6] GHCC: https://github.com/huzecong/ghcc. With the help of a decompiler, each binary is lifted into equivalent C pseudocode and processed into a list of lexemes and an implementation-defined format for enumerating source-level variable information. Because the DIRE dataset is a preprocessed dataset that does not conform with ContraBin’s expected inputs, we instead obtained the list of GitHub repositories that were used to construct the DIRE dataset and compiled each project ourselves. We modified GHCC to save the intermediate LLVM bitcode objects produced during compilation and disassemble each bitcode object into readable LLVM assembly using the LLVM disassembler.
@@
In our experiments, we select function names with number of functions larger than 200, and remove some function names that are not specific to a type of function (i.e., main, _, and __list_add) to make sure the model is trained on meaningful data. We further strip all LLVM files of their original function names (i.e., @TESTFUN0 as the name of the first function in an LLVM file) and measure ContraBin’s ability to classify these function names. The dataset statistics are shown in Table 2. We fine-tune each model with 5 epochs and a block size of 256. We use training and validation batch sizes of 8 and 16, respectively, and choose learning rate to be 2e-5 and maximal gradient normalization to be 1. In all fine-tuning processes, we use the default random seed as 123456.
@@
AnghaBench For the binary code summarization and reverse engineering tasks, we use the AnghaBench test set (Da Silva et al., 2021) during pre-training, which includes code never been seen by the model. Specifically, we only select the main function in LLVM code and its paired source code for fine-tuning the code summarization and reverse engineering tasks. To better fit the capacity of pre-trained language models, we further truncate the length of binary code by only selecting the first 512 instructions. The dataset statistics are displayed in Table 2. In this paper, we evaluate the model’s ability to translate and summarize binary code using the validation set and analyze its embedding using the test set.
Table 2: Dataset statistics for algorithmic classification.